NPP Projection Close 8/7
To do this evening/ was completed:
Presentation Issues: Some guests complained about the volume being too low for Weapons in theater 3 this evening. I raised it from 5.5 to 6.5 and, after getting a chance to check the room myself, ended up brining it to 7.0 since the overall mix is generally very quiet and there are many instances throughout the film where the sound drops out entirely to build tension (I happened to watch it this morning at NWB and texted Joe asking to raise the volume since it sounded low with the handful of staff there -- I have raised the feature volume on all playlists to avoid any further complaints)
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Setup for Spoons Toons & Booze this evening was very quick thanks to Alec's organizing the necessary equipment this afternoon...the hosts had the show started promptly and all went smoothly (tidied up the gear and placed it out of the way as the STB hosts will be back on Saturday for a brunch show -- mics are back in the booth to charge)
Ordering/Receiving (posters & trailers): Additional posters for HIM were delivered this afternoon and stored behind 3/7...still waiting on the Realm of Satan one sheets to arrive
Ordering/Receiving (content): Began converting Realm of Satan to DCP on the 4/6 desktop and saved the file to the George Ivan drive
Presentation Issues: Some guests complained about the volume being too low for Weapons in theater 3 this evening. I raised it from 5.5 to 6.5 and, after getting a chance to check the room myself, ended up brining it to 7.0 since the overall mix is generally very quiet and there are many instances throughout the film where the sound drops out entirely to build tension (I happened to watch it this morning at NWB and texted Joe asking to raise the volume since it sounded low with the handful of staff there -- I have raised the feature volume on all playlists to avoid any further complaints)
Preshow Clips/Playlist:
Trailer Updates:
Special Events/ AV set up: Setup for Spoons Toons & Booze this evening was very quick thanks to Alec's organizing the necessary equipment this afternoon...the hosts had the show started promptly and all went smoothly (tidied up the gear and placed it out of the way as the STB hosts will be back on Saturday for a brunch show -- mics are back in the booth to charge)
Ordering/Receiving (posters & trailers): Additional posters for HIM were delivered this afternoon and stored behind 3/7...still waiting on the Realm of Satan one sheets to arrive
Ordering/Receiving (content): Began converting Realm of Satan to DCP on the 4/6 desktop and saved the file to the George Ivan drive
David Jennings sent us DCP content for Bad Boys through Deluxe in the event that we are unable to run the print given the missing reel (see below)
Outgoing/Shipping (content):
Print/video Inspection: Spent the majority of the evening inspecting Bad Boys which, unfortunately, is missing its second reel but, despite this, is in good condition overall. An inspection report was filled out and shared with David Jennings, who I contacted and discussed some workarounds to this (a possible replacement print and/or last ditch DCP file which was sent out via Deluxe this evening but is not yet visible on the portal). I took some stills of the print for posterity's sake and shared those with Conrad and Desmond. Will update as the situation develops and we hear back from David/Sony regarding the print
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Double checked tomorrow's cues and schedule and initialed for confirmation
Marquee updated?: Removed the 35mm titles and added letters for the marquee rental for tomorrow afternoon (the letters previously up have been organized and placed on the workbench behind 3/7)
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Outgoing/Shipping (content):
Print/video Inspection: Spent the majority of the evening inspecting Bad Boys which, unfortunately, is missing its second reel but, despite this, is in good condition overall. An inspection report was filled out and shared with David Jennings, who I contacted and discussed some workarounds to this (a possible replacement print and/or last ditch DCP file which was sent out via Deluxe this evening but is not yet visible on the portal). I took some stills of the print for posterity's sake and shared those with Conrad and Desmond. Will update as the situation develops and we hear back from David/Sony regarding the print
Asset tracking doc updated/checked?:
Maintenance/cleaning:
Cue sheets updated/checked (every shift): Double checked tomorrow's cues and schedule and initialed for confirmation
Marquee updated?: Removed the 35mm titles and added letters for the marquee rental for tomorrow afternoon (the letters previously up have been organized and placed on the workbench behind 3/7)
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment